# Primordial Black Hole Formation via Inverted Bubble Collapse

### v1

@@ Abstract
We propose a novel mechanism of primordial black hole (PBH) formation through inverted bubble collapse. In this scenario, bubbles nucleate sparsely in an incomplete first-order phase transition, followed by a bulk phase transition in the rest of the universe that inverts these pre-existing bubbles into false vacuum regions. These spherically symmetric false-vacuum bubbles subsequently collapse to form PBHs. Unlike conventional PBH formation mechanisms associated with domain wall collapse or bubble coalescence, our inverted bubble collapse mechanism naturally ensures spherical collapse. We demonstrate that, when applied to the electroweak phase transition, this mechanism can produce highly monochromatic PBHs with masses up to ${\cal O}(10^{-6}\,\text{--}\,10^{-5})M_{\odot}$, which potentially explain the microlensing events observed in the OGLE and Subaru HSC data.
@@ I Introduction
The composition of our Universe presents one of the most profound mysteries in modern cosmology. Observational evidence indicates that approximately one-quarter of the energy density of the Universe consists of dark matter, whose nature remains unknown Aghanim et al. (2020); Rubin and Ford (1970); Zwicky (1933). Among various candidates proposed to explain dark matter, primordial black holes (PBHs) stand out as an intriguing possibility Hawking (1972); Carr and Hawking (1974); Carr (1975); Frampton et al. (2010); Carr et al. (2016). While PBHs could potentially account for all of the dark matter, they might also coexist with other dark matter components. Indeed, scenarios in which PBHs constitute a non-negligible fraction or even all of the dark matter have been widely studied. Carr et al. (2021); Carr and Kuhnel (2020); Green and Kavanagh (2021); Frampton et al. (2010).
@@
The formation of PBHs requires significant density perturbations in the early Universe. Many proposed mechanisms rely on superhorizon-scale fluctuations, with one widely studied scenario involving the generation of large curvature perturbations during inflation that later reenter the horizon and collapse into PBHs Zel’dovich and Novikov (1967); Hawking (1972); Carr and Hawking (1974); Carr (1975); Yokoyama (1997); García-Bellido et al. (1996); Kawasaki et al. (1998). However, such scenarios often require fine-tuning of the inflaton potential, particularly in creating ultra-flat regions in the case of single-field inflation models, which imposes non-trivial constraints on inflationary model building. Another possibility is to make use of the large isocurvature perturbations of baryons or dark matter. For instance, PBH formation from collapsing baryon/axion bubbles represents an interesting realization of such a mechanism Dolgov and Silk (1993); Dolgov et al. (2009); Kawasaki and Murai (2019); Kitajima and Takahashi (2020); Kawasaki et al. (2021); Kasai et al. (2022, 2023, 2024a, 2024b).
@@
Alternative mechanisms for generating large overdensities include those involving topological defects such as domain walls Ferrer et al. (2019); Ge (2020); Liu et al. (2020); Gelmini et al. (2023); Kitajima et al. (2024); Ge et al. (2024); Gouttenoire and Vitagliano (2024a, b); Ferreira et al. (2024); Dunsky and Kongsore (2024); Gouttenoire et al. (2025) and first-order phase transitions Sato et al. (1981); Kodama et al. (1981, 1982); Maeda et al. (1982); Jedamzik and Niemeyer (1999); Liu et al. (2022); Hashino et al. (2022, 2023); Kawana et al. (2023); Lewicki et al. (2023); Gouttenoire and Volansky (2024); Flores et al. (2024); Hashino et al. (2025). However, these scenarios face fundamental challenges, as the overdense regions are expected to deviate significantly from spherical symmetry, and it remains unclear how much deviation is allowed for successful PBH formation Yoo et al. (2020). Furthermore, even if we restrict ourselves to cases where overdense regions are approximately spherically symmetric, it is difficult to estimate the probability of forming such nearly spherical overdense regions. These uncertainties make it challenging to predict the PBH abundance and understand the collapse process.
@@
In this Letter, we propose a novel mechanism for PBH formation through inverted bubble collapse in an incomplete first-order phase transition followed by a bulk transition. In our scenario, bubbles sparsely nucleate during an incomplete first-order phase transition. These bubbles are spherically symmetric when nucleated Coleman et al. (1978), and maintain this symmetry unless they collide. Crucially, before this transition is completed, a second phase transition occurs throughout the Universe except within the bubbles, transforming the bubble interiors into false vacuum regions. Consequently, the initially expanding bubbles begin to contract and eventually collapse. When sufficiently energetic bubbles collapse, they can form PBHs while naturally preserving spherical symmetry. Unlike PBH formation scenarios involving topological defect collapse or bubble coalescence, our mechanism allows for precise prediction of PBH abundance in terms of model parameters due to this preserved spherical symmetry.
@@
While this mechanism itself is a general PBH formation scenario not limited to any specific phase transitions, we demonstrate its realization in the context of the electroweak phase transition (EWPT) by extending the Standard Model (SM) with a singlet scalar field. In this setup, an initial (incomplete) first-order phase transition occurs in the singlet direction, followed by the EWPT in the Higgs direction. We show that this scenario can produce highly monochromatic PBHs with masses up to ${\cal O}(10^{-6}\,\text{--}\,10^{-5})M_{\odot}$, which could potentially explain the microlensing events observed in the OGLE and Subaru HSC data Mroz et al. (2017); Niikura et al. (2019a, b) and will also be probed by the ongoing Subaru HSC searches for PBHs and other compact objects. Depending on the model parameters, it can also generate PBHs with much smaller masses that could explain all dark matter. Furthermore, if the EWPT is also of first-order, it could generate gravitational waves detectable by future space-borne interferometers such as LISA Arun et al. (2022) (see also Refs. Hashino et al. (2017); Vaskonen (2017); Beniwal et al. (2017); Hashino et al. (2019); Ellis et al. (2019, 2019); Alves et al. (2019); Alanne et al. (2020); Ellis et al. (2023); Ramsey-Musolf et al. (2024) for the studies of theoretical predictions of gravitational waves in the extension of the SM with a real singlet scalar.) Indeed, as we will see, for the model parameters we adopt, the phase transition in the Higgs direction is of first order.
II PBH formation in the inverted bubble collapse
@@
In this section, we derive the mass distribution of PBHs formed through our inverted bubble collapse mechanism. We first calculate the bubble size distribution, then determine their collapse conditions and derive the mass spectrum of PBHs.
@@
To calculate the bubble size distribution, we first need to track the evolution of bubbles in the cooling universe. We evaluate this at time $t=t_{\mathrm{c}}$, when the bubble walls start to shrink after their interiors have become false vacuum regions through the bulk phase transition. Note that $t_{c}$ is generically later than the bulk phase transition that occurs at $t=t_{\mathrm{PT}}$. The bubble nucleation rate per unit volume, $\Gamma(T)$, depends on the cosmic temperature $T$ and grows exponentially as the universe cools down.
@@
$\frac{\mathrm{d}n_{\mathrm{b}}}{\mathrm{d}t}=-3Hn_{\mathrm{b}}+\Gamma(T(t))\ .$ (1)
The solution to this equation is
@@
where $a$ is the scale factor, and $R_{\mathrm{in}}$ is the physical critical radius at the bubble nucleation. In the second equality, we used $a\propto t^{1/2}$, as we focus on PBH formation during the radiation-dominated (RD) era in this Letter. As a result, the size distribution of the bubble at $t=t_{c}$ is given by
@@
$\frac{\mathrm{d}n_{\mathrm{b}}}{\mathrm{d}R}(t_{c})$ $\simeq\frac{\Gamma(T(t_{R}))}{v}\frac{t_{R}^{2}}{t_{c}^{2}}$
$\simeq\frac{\Gamma(T(t_{R}))}{v}\left(1-\frac{R}{2vt_{c}}\right)^{4}\ .$ (6)
@@
Next, we discuss the PBH formation from a false bubble. Here, we denote the potential energy difference between the inside and outside of the bubble by $\Delta V$, which we treat as constant during the relevant timescale of bubble collapse. In realistic situations, $\Delta V$ depends on temperature, as we will see later; however, for simplicity, we treat it as a constant in the following analysis.
Let us assume that the radius of the bubble is $R$ when its interior transitions to the false vacuum and the bubble wall starts to shrink. Then, the energy of the bubble can be approximately expressed as
@@
As the bubble shrinks, a part of the vacuum energy swept by the bubble wall is accumulated in the kinetic energy of the wall, and this energy gets progressively concentrated inside a smaller region. When the first-order phase transition is sufficiently strong with a large vacuum energy difference, the bubble wall can undergo run-away acceleration despite the dissipation through interactions with the ambient plasma Bodeker and Moore (2009). In such run-away cases, we introduce an efficiency parameter $\epsilon(\leq 1)$ to characterize the fraction of vacuum energy converted into the wall’s kinetic energy. Then, if the bubble size becomes smaller than the Schwarzschild radius, $R_{\mathrm{s}}=2G\epsilon M_{\mathrm{b}}$, the bubble collapses into a PBH. Considering that the bubble can be as small as the width of the bubble wall, $\delta$, the condition for the PBH formation is given by [^1] In fact, the bubble wall width becomes thinner due to the Lorentz contraction as it gets accelerated. The wall width $\delta$ is assumed to include this effect.
@@
Considering the upper bound of the bubble radius, we obtain the allowed range of the PBH mass as
@@
where $H_{c}$ and $T_{c}$ denote the Hubble parameter and temperature when $t=t_{c}$, respectively. In the following, we assume that $M_{\mathrm{min}}$ is much smaller than the mass region of interest, which is a condition typically satisfied in realistic situations. [^2] The width of the bubble wall in the rest frame is typically estimated by the inverse of the mass scale of the field driving the phase transition. For instance, if we take the mass scale to be 100 GeV, we obtain $M_{\mathrm{min}}\simeq 6\times 10^{-22}M_{\odot}/\gamma$ with $\gamma$ being the Lorentz factor of the wall.
@@
Next, we evaluate the PBH mass function assuming the temperature dependence of the bubble nucleation rate. The bubble nucleation rate is considered to increase progressively toward the bulk phase transition that occurs at temperature $T=T_{\rm PT}(\geq T_{c})$, and then effectively drops down to zero afterwards. Thus, around the bulk phase transition, it can be parametrized as [^3] We require that the first-order phase transition via bubble nucleation should not occur inside the bubbles, as it would spoil the spherical symmetry and reduce the energy inside them. This condition is met in the case of the singlet extension of the SM studied in the next section.
$\Gamma(T)\simeq\begin{cases}cH_{\mathrm{PT}}^{4}\,\,e^{-\alpha\frac{T-T_{\mathrm{PT}}}{T_{\mathrm{PT}}}}&\mathrm{~{}~{}for~{}~{}}T\geq T_{\mathrm{PT}}\\ 0&\mathrm{~{}~{}for~{}~{}}T\leq T_{\mathrm{PT}}\end{cases},$ (13)
@@
$T_{R}$ $=\sqrt{\frac{t_{c}}{t_{R}}}T_{c}$
$\simeq\left(1-\frac{R}{2vt_{c}}\right)^{-1}T_{c}$
@@
Then, the mass function is given by
$\frac{\mathrm{d}\rho_{\mathrm{PBH}}}{\mathrm{d}\ln M}(t)$
$\simeq\frac{cMH_{\mathrm{PT}}^{4}}{3v}\left(\frac{3M}{4\pi\epsilon\Delta V}\right)^{1/3}\left[1-\frac{H_{c}}{v}\left(\frac{3M}{4\pi\epsilon\Delta V}\right)^{1/3}\right]^{4}\left(\frac{a(t_{c})}{a(t)}\right)^{3}$
×$~{}~{}~{}~{}\times\exp\left[-\alpha\left(\frac{T_{c}}{T_{\mathrm{PT}}}\left[1-\frac{H_{c}}{v}\left(\frac{3M}{4\pi\epsilon\Delta V}\right)^{1/3}\right]^{-1}-1\right)\right]\ ,$ (15)
for $M$ satisfying $M_{\mathrm{min}}<M<M_{\mathrm{max}}$ and $T_{R}>T_{\mathrm{PT}}$. This mass function is translated to the energy ratio to dark matter by
$\frac{\mathrm{d}f_{\mathrm{PBH}}}{\mathrm{d}\ln M}=$ $\frac{1}{\rho_{\mathrm{DM}}(t)}\frac{\mathrm{d}\rho_{\mathrm{PBH}}}{\mathrm{d}\ln M}(t)$
$\simeq$ $\frac{1}{0.44\,\mathrm{eV}}\frac{45}{2\pi^{2}g_{*s}(T_{c})T_{c}^{3}}\frac{cMH_{\mathrm{PT}}^{4}}{3v}\left(\frac{3M}{4\pi\epsilon\Delta V}\right)^{1/3}$
×$\times\left[1-\frac{H_{c}}{v}\left(\frac{3M}{4\pi\epsilon\Delta V}\right)^{1/3}\right]^{4}$
@@
In Fig. 1, we show the predicted PBH abundance in this model. We use the parameter sets (a) and (b) shown in Tab. 1 for the red and blue solid lines, respectively. These parameter sets correspond to a constant bubble nucleation rate ($\alpha=0$), fast bubble expansion ($v=1$), and efficient accumulation of the false vacuum energy ($\epsilon=1$). For comparison, we also show the results for the cases with $\alpha=10$ (red dashed line) and with $v=0.5$ (red dot-dashed line) based on the parameter set (a). Note that $v$, $\epsilon$, and $\Delta V$ appear in Eq. (16) in the form of $v(\epsilon\Delta V)^{1/3}$, and the changes of their values are degenerate in $\mathrm{d}f_{\mathrm{PBH}}/\mathrm{d}\ln M$. In these cases, the mass distribution follows $\mathrm{d}f_{\mathrm{PBH}}/\mathrm{d}\ln M\propto M^{4/3}$ for $M\ll M_{\mathrm{max}}$ and has a cutoff at $M=M_{\mathrm{max}}$. On the other hand, the PBH mass function for the parameter set (b) has a cutoff on smaller mass. This is because all bubbles expand at least between $t_{\mathrm{PT}}$ and $t_{c}$, and there is a minimal bubble radius at $t=t_{c}$, leading to the lower bound on the PBH mass. As we will see below, the PBH mass distribution approaches a monochromatic distribution as $t_{c}$ is significantly delayed from $t_{\mathrm{PT}}$. If we set $T_{\mathrm{PT}}$ equal to $T_{c}$ in the parameter set (b) and tune $c$, we obtain the same mass function, except that there is no cutoff for small masses as shown in the blue dashed line.

### v3

@@ Abstract
We propose a novel mechanism of primordial black hole (PBH) formation through inverted bubble collapse. In this scenario, bubbles nucleate sparsely in an incomplete first-order phase transition, such that they remain isolated and do not percolate or collide with each other due to the extremely low nucleation rate. This is followed by a bulk phase transition in the rest of the universe that inverts these pre-existing bubbles into false vacuum regions. These spherically symmetric false-vacuum bubbles subsequently collapse to form PBHs. Unlike conventional PBH formation mechanisms associated with domain wall collapse or bubble coalescence, our inverted bubble collapse mechanism naturally ensures spherical collapse. We demonstrate that, when applied to the singlet extension of the Standard Model, this mechanism can produce highly monochromatic PBHs with masses up to ${\cal O}(10^{-7}\,\text{--}\,10^{-5})M_{\odot}$, which potentially explain the microlensing events observed in the OGLE and Subaru HSC data.
@@ I Introduction
The composition of our universe presents one of the most profound mysteries in modern cosmology. Observational evidence indicates that approximately one-quarter of the energy density of the universe consists of dark matter, whose nature remains unknown [1, 2, 3]. Among various candidates proposed to explain dark matter, primordial black holes (PBHs) stand out as an intriguing possibility [4, 5, 6, 7, 8]. While PBHs could potentially account for all of the dark matter, they might also coexist with other dark matter components. Indeed, scenarios in which PBHs constitute a non-negligible fraction or even all of the dark matter have been widely studied. [9, 10, 11, 7].
@@
The formation of PBHs requires significant density perturbations in the early universe. Many proposed mechanisms rely on superhorizon-scale fluctuations, with one widely studied scenario involving the generation of large curvature perturbations during inflation that later reenter the horizon and collapse into PBHs [12, 4, 5, 6, 13, 14, 15]. However, such scenarios often require fine-tuning of the inflaton potential, particularly in creating ultra-flat regions in the case of single-field inflation models, which imposes non-trivial constraints on inflationary model building. Another possibility is to make use of the large isocurvature perturbations of baryons or dark matter. For instance, PBH formation from collapsing baryon/axion bubbles represents an interesting realization of such a mechanism [16, 17, 18, 19, 20, 21, 22, 23, 24].
@@
Alternative mechanisms for generating large overdensities include those involving topological defects such as domain walls [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35] and first-order phase transitions (FOPTs) [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. However, these scenarios face fundamental challenges, as the overdense regions are expected to deviate significantly from spherical symmetry, and it remains unclear how much deviation is allowed for successful PBH formation (see Refs. [55, 56, 57] for the discussion on the PBH formation from Gaussian curvature perturbations). These uncertainties make it challenging to predict the PBH abundance and understand the collapse process.
@@
In this paper, we propose a novel mechanism for PBH formation through inverted bubble collapse (IBC) in an incomplete FOPT followed by a bulk phase transition. In our scenario, the FOPT is incomplete because the bubble nucleation rate is so low that bubbles remain isolated and the transition never completes through percolation. These bubbles are spherically symmetric when nucleated [58], and maintain this symmetry unless they collide. Crucially, before this FOPT is completed, a second phase transition occurs throughout the universe except within the bubbles, transforming the bubble interiors into false vacuum regions. Consequently, the initially expanding bubbles begin to contract and eventually collapse. When sufficiently energetic bubbles collapse, they can form PBHs while naturally preserving spherical symmetry. Unlike PBH formation scenarios involving topological defect collapse or bubble coalescence, our mechanism allows for precise prediction of PBH abundance in terms of model parameters thanks to this preserved spherical symmetry.
@@
We demonstrate that our mechanism can be implemented in the context of the electroweak phase transition (EWPT) by extending the Standard Model (SM) with singlet scalar fields. In this setup, an initial incomplete FOPT and the subsequent bulk phase transition occur in the singlet sector, followed by the EWPT in the Higgs direction. We show that this scenario can produce highly monochromatic PBHs with masses up to ${\cal O}(10^{-7}\,\text{--}\,10^{-5})M_{\odot}$, which could potentially explain the microlensing events observed in the OGLE and Subaru HSC data [59, 60, 61] and will also be probed by the ongoing Subaru HSC searches for PBHs and other compact objects. Depending on the model parameters, it can also generate PBHs with much smaller masses that could explain all dark matter. Furthermore, if either the bulk phase transition in the singlet sector or the EWPT is of first-order, it could generate gravitational waves detectable by future space-borne interferometers such as LISA [62] (see also Refs. [63, 64, 65, 66, 67, 67, 68, 69, 70, 71] for theoretical studies of gravitational waves in the singlet extension).
Lastly, we note that the expanding and shrinking bubbles have also been studied in Ref. [72], which focused on the bubble nucleation during the very early reheating stage. In their scenario, the shrinking bubbles leave small density perturbations that later collapse to PBHs, which grow via mass accretion during the inflaton matter dominated universe. In contrast, our mechanism operates in the radiation dominated (RD) universe and forms PBHs directly through bubble collapse. Furthermore, our approach allows us to calculate the PBH mass spectrum solely from the phase transition parameters, independent of specific inflation models and reheating processes.
II PBH formation in the IBC mechanism
@@
In this section, we derive the mass distribution of PBHs formed through our IBC mechanism. We first describe the outline of our IBC mechanism. Then, we calculate the bubble size distribution, determine their collapse conditions, and derive the mass spectrum of PBHs.
II.1 Outline of the IBC mechanism
In our IBC mechanism, an incomplete FOPT is followed by a bulk phase transition at $T=T_{\mathrm{PT}}$. At high temperatures, the FOPT proceeds via bubble nucleation. As long as the nucleation rate per unit volume, $\Gamma(T)$, satisfies $\Gamma(T)\ll H^{4}$, where $H$ is the Hubble parameter, bubbles of the true vacuum are formed only in a tiny fraction of Hubble volumes, and the FOPT hardly progresses even though each bubble expands with velocity $v$. Once $\Gamma$ becomes comparable to or larger than $H^{4}$, bubbles percolate, and the FOPT would normally complete. In our scenario, we instead consider the situation where the bulk phase transition occurs while $\Gamma(T)\ll H^{4}$, so the FOPT remains incomplete, leaving isolated bubbles with a number density below one bubble per Hubble volume at $T=T_{\mathrm{PT}}$. After the bulk phase transition, $\Gamma(T)$ becomes zero or negligibly small, and the bubble nucleation ceases. As the universe cools down, the bulk region transitions to the true vacuum, while the bubble interiors effectively become regions of false vacuum. Then, the bubbles start to shrink at $T=T_{c}$. If the energy of the false vacuum is efficiently converted into the kinetic energy of the bubble wall, and the total energy is concentrated within the Schwarzschild radius, the contracting bubbles collapse into PBHs. A schematic illustration of the IBC mechanism is shown in Fig. 1.
Figure 1: Schematic illustration of the IBC mechanism. Left panel: At high temperatures, bubbles nucleate at a low rate, and isolated bubbles expand with velocity $v$. Center panel: After the bulk phase transition at $T=T_{\mathrm{PT}}$, bubble nucleation ceases, while the bubbles continue to expand. Right panel: For $T<T_{c}$, the bubbles shrink and collapse into PBHs.
II.2 Bubble size distribution
@@
To calculate the bubble size distribution, we first need to track the evolution of bubbles in the cooling universe. We evaluate this at time $t=t_{c}$ corresponding to $T=T_{c}$, when the bubble walls start to shrink after their interiors have become false vacuum regions through the bulk phase transition. Note that $t_{c}$ is generically later than the bulk phase transition that occurs at $t=t_{\mathrm{PT}}$ corresponding to $T=T_{\mathrm{PT}}$.
@@
$\frac{\mathrm{d}n_{\mathrm{b}}}{\mathrm{d}t}=-3Hn_{\mathrm{b}}+\Gamma(T(t))\ ,$ (1)
where $\Gamma(T)$ is the bubble nucleation rate per unit volume, which depends on the cosmic temperature $T$ and grows exponentially as the universe cools down. The solution to this equation is
@@
where $a$ is the scale factor, and $R_{\mathrm{in}}$ is the physical critical radius at the bubble nucleation. In the second equality, we used $a\propto t^{1/2}$, as we focus on PBH formation during the RD era in this paper. As a result, the size distribution of the bubble at $t=t_{c}$ is given by
@@
$\frac{\mathrm{d}n_{\mathrm{b}}}{\mathrm{d}R}(t_{c})$ $\simeq\frac{\Gamma(T(t_{R}))}{v}\left(1-\frac{R}{2vt_{c}}\right)^{4}\ .$ (6)
@@
II.3 PBH formation
Next, we discuss the PBH formation from a false vacuum bubble. Here, we denote the potential energy difference between the inside and outside of the bubble by $\Delta V$, which we treat as constant during the relevant timescale of bubble collapse. In realistic situations, $\Delta V$ depends on temperature, as we will see later; however, for simplicity, we treat it as a constant in the following analysis.
Let us assume that the radius of the bubble is $R$ when the bubble wall starts to shrink after its interior transitions to the false vacuum. Then, the energy of the bubble can be approximately expressed as
@@ II.3 PBH formation
Under certain situations, when the energy difference is sufficiently large, the bubble wall can undergo runaway acceleration even in cases with plasma dissipation, whereas in the absence of dissipation, runaway is naturally achieved [73, 74]. Then, as the bubble shrinks, a part of the vacuum energy swept by the bubble wall is accumulated in the kinetic energy of the wall, and this energy gets progressively concentrated inside a smaller region. In such runaway cases, we introduce an efficiency parameter $\epsilon(\leq 1)$ to characterize the fraction of vacuum energy converted into the wall’s kinetic energy. Then, if the bubble size becomes smaller than the Schwarzschild radius, $R_{\mathrm{s}}=2G\epsilon M_{\mathrm{b}}$, the bubble collapses into a PBH. Considering that the bubble can be as small as the width of the bubble wall, $\delta$, the condition for the PBH formation is given by [^1] In fact, the bubble wall width becomes thinner due to the Lorentz contraction as it gets accelerated. The wall width $\delta$ here is assumed to include this effect.
@@
Considering the upper bound of the bubble radius, we obtain the possible range of the PBH mass as
@@
where $H_{c}$ denotes the Hubble parameter when $t=t_{c}$. In the following, we assume that $M_{\mathrm{min}}$ is much smaller than the mass region of interest, which is a condition typically satisfied in realistic situations. [^2] The width of the bubble wall in the rest frame is typically estimated by the inverse of the mass scale of the field driving the phase transition. For instance, if we take the mass scale to be 100 GeV, we obtain $M_{\mathrm{min}}\simeq 6\times 10^{-22}M_{\odot}/\gamma$ with $\gamma$ being the Lorentz factor of the wall.
II.4 PBH mass function
@@
Next, we evaluate the PBH mass function assuming the temperature dependence of the bubble nucleation rate. The bubble nucleation rate is considered to increase progressively toward the bulk phase transition at $T=T_{\rm PT}(\geq T_{c})$, and then effectively drops down to zero afterwards. Thus, around the bulk phase transition, it can be parametrized as [^3] We require that the FOPT via bubble nucleation should not occur inside the bubbles, as it would spoil the spherical symmetry and reduce the energy inside them. This condition is met in the case of the singlet extension of the SM studied in the next section.
$\Gamma(T)\;\simeq\;cH_{\mathrm{PT}}^{4}\,\,e^{-\alpha\frac{T-T_{\mathrm{PT}}}{T_{\mathrm{PT}}}}$ $\mathrm{~{}~{}for~{}~{}}T\geq T_{\mathrm{PT}},$ (13)
@@
$T_{R}$ $=\sqrt{\frac{t_{c}}{t_{R}}}T_{c}\simeq\left(1-\frac{R}{2vt_{c}}\right)^{-1}T_{c}$
@@
Then, we obtain the energy ratio to dark matter by
$\frac{\mathrm{d}f_{\mathrm{PBH}}}{\mathrm{d}\ln M}\equiv$ $\frac{1}{\rho_{\mathrm{DM}}(t)}\frac{\mathrm{d}\rho_{\mathrm{PBH}}}{\mathrm{d}\ln M}(t)$
$\simeq$ $\frac{1}{0.44\,\mathrm{eV}}\frac{45}{2\pi^{2}g_{*s}(T_{c})T_{c}^{3}}\frac{cMH_{\mathrm{PT}}^{4}}{3v}\left(\frac{3M}{4\pi\epsilon\Delta V}\right)^{1/3}\left[1-\frac{H_{c}}{v}\left(\frac{3M}{4\pi\epsilon\Delta V}\right)^{1/3}\right]^{4}$
@@
for $M$ satisfying $M_{\mathrm{min}}<M<M_{\mathrm{max}}$ and $T_{R}>T_{\mathrm{PT}}$. Here, we used $\rho_{\mathrm{DM}}/s=0.44$ eV with $s$ being the entropy density.
In Fig. 2, we show the predicted PBH abundance in our scenario. We use the parameter sets (a) and (b) shown in Tab. 1 for the red and blue solid lines, respectively. These parameter sets correspond to a constant bubble nucleation rate ($\alpha=0$), fast bubble expansion ($v=1$), and efficient accumulation of the false vacuum energy ($\epsilon=1$). For comparison, we also show the results for the cases with $\alpha=10$ (red dashed line) and with $v=0.5$ (red dot-dashed line) based on the parameter set (a). Note that $v$, $\epsilon$, and $\Delta V$ appear in Eq. (15) in the form of $v(\epsilon\Delta V)^{1/3}$, and the changes of their values are degenerate in $\mathrm{d}f_{\mathrm{PBH}}/\mathrm{d}\ln M$. In these cases, the mass distribution follows $\mathrm{d}f_{\mathrm{PBH}}/\mathrm{d}\ln M\propto M^{4/3}$ for $M\ll M_{\mathrm{max}}$ and has a cutoff at $M=M_{\mathrm{max}}$. On the other hand, the PBH mass function for the parameter set (b) has a cutoff on smaller mass. This is because all the bubbles continue to expand at least between $t_{\mathrm{PT}}$ and $t_{c}$, and there is a minimal bubble radius at $t=t_{c}$, leading to the lower bound on the PBH mass. As we will see below, the PBH mass distribution approaches a monochromatic distribution as $t_{c}$ is significantly delayed from $t_{\mathrm{PT}}$. Setting $T_{\mathrm{PT}}=T_{c}$ in parameter set (b) yields the same mass function for a specific choice of $c$, but without the cutoff for small masses (blue dashed line).